OpenStudy (anonymous):

In October, a hardware store purchased snow shovels for $8 each. The original markup was 50% based on the selling price. In December, the store had a sale and marked down the shovels by 20%. By January 1, the sale was over and the shovels were marked up 15%. In March, the store held a clearance sale of all winter tools by having a final markdown of 60%. What was the final selling price of the shovels?

OpenStudy (mertsj):

1. $8 + 50% of $8 = original selling price 2. December sale price: 80% of original selling price 3. January 1 price: 115% of December sale price 4. March clearance sale price: 40% of January 1 price.

OpenStudy (anonymous):

Im sorry but i still didnt get it .. Pls need help more?

OpenStudy (mertsj):

So you can't do step 1 or what?

OpenStudy (anonymous):

No mean do i have to add the percentage after adding 8+50%= the upcoming month as well?

OpenStudy (mertsj):

What is 50% of 8?

OpenStudy (mertsj):

Ok. Let me rephrase that. What is half of 8 ?

OpenStudy (anonymous):

4

OpenStudy (mertsj):

Add that number to 8. 4+8= ???

OpenStudy (anonymous):

12

OpenStudy (mertsj):

1. $8 + 50% of $8 = original selling price=12

OpenStudy (mertsj):

So now we have the answer to Step 1

OpenStudy (mertsj):

2. December sale price: 80% of original selling price Now we know that the original selling price was $12 and we can answer Step 2 by finding 80% of the original selling price. What is 80% of 12?

OpenStudy (anonymous):

Oh ok i see

OpenStudy (anonymous):

9.6

OpenStudy (mertsj):

So now we know that the December sale price is $9.60 and we can answer Step 3

OpenStudy (mertsj):

3. January 1 price: 115% of December sale price What is 115% of $9.60 ???

OpenStudy (anonymous):

11.04

OpenStudy (mertsj):

So now we know that the January 1 price was $11.04 and we can do step 4 4. March clearance sale price: 40% of January 1 price. What is 40% pf $11.04 ???

OpenStudy (anonymous):

4.41

OpenStudy (mertsj):

I believe that 4.416 would round to 4.42

OpenStudy (mertsj):

So $4.42 is the final clearance sale price.

OpenStudy (anonymous):

Yes it ..Ok so we round it great

OpenStudy (anonymous):

Oh my god thank you so much i will write this down thank you thank you

OpenStudy (mertsj):

yw